OpenStudy (anonymous):

What is the equation of the line passing through (8, –6) and perpendicular to the line y = x – 2 in slope-intercept form?

OpenStudy (anonymous):

use point slope form. y - y1 = m (x - x1) find the slope. remember, if it's perpendicular, it's the negative reciprocal. in this case it's a slope of -1 plug in all your values. I'll get you started y1 = -6 x1 = 8 m = -1 y - y1 = m (x - x1)

OpenStudy (radar):

Can you work it now?

OpenStudy (anonymous):

imma tried to do my best....

OpenStudy (radar):

Plug in the values in the formula that ZhangYan gave you like this: y-(-6)=-1(x-8)

OpenStudy (anonymous):

Thank you so very much!!

OpenStudy (radar):

You're welcome, the results could look like this: y+6=-x+8 y=-x+2